OpenStudy (anonymous):

f(x)= 2/x-5 and g(x)= 2/x +5

OpenStudy (anonymous):

whats the question here?

OpenStudy (shamil98):

what's the question?

OpenStudy (anonymous):

how do I solve for (fog)

OpenStudy (shamil98):

( f o g ) means you substitute the equation of g(x) into f(x) f(g(x)) = 2/(2/x+5) -5

OpenStudy (shamil98):

then simplify.

OpenStudy (anonymous):

well I'm stuck at that point

OpenStudy (shamil98):

\[f(g(x)) = 2 \times \frac{ x+5 }{ 2 } - 5\]

OpenStudy (anonymous):

do i multiply x/2 for the top and bottom ???

OpenStudy (shamil98):

f(g(x)) = x

OpenStudy (shamil98):

when dividng fractions you multiply the recipocral

OpenStudy (shamil98):